OpenStudy (anonymous):

Ok, I am so not liking these problems. Make: 1. " a " the subject of "d = sqrt(a) / n (square root of "a" OVER "n") 2. "l" the subject of "T = 1/5 (sqrt"l")" [that is one fifth to the square root of "l"] 3. "a" the subject of "c = sqrt a^2-b^2 " Its a bit complicated to understand. Just respond if any probs.

jhonyy9 (jhonyy9):

1 is d= radical a /n ? 2 T=1/5radical i ? 3 c = radical a*2 -b*2 ???

jhonyy9 (jhonyy9):

or 1. d=a*2 /n 2 T=I*2 /5 3. ???

OpenStudy (anonymous):

I thought it would be like "a = ____" or "l = ____" since it says to make it the SUBJECT OF. U get me?

OpenStudy (anonymous):

But I do thank you for at least attempting to find an answer. :)